\begin{thebibliography}{99}


%==================
\bibitem{Shannon:1948}
C.~Shannon, \emph{A mathematical theory of communication},  Bell System Techn. J.,
\textbf{27} (1948),  3, 379--423; \textbf{27} (1948),  4, 623--656.

%%==================
%\bibitem{petermann}  A. Splett, C. Kurzke, and K. Petermann,
%In Proc. of the European Conf. on Opt. Com., ECOC’93, \textbf{2}, 41 (1993).

%==================
\bibitem{Mitra:2001}
P. P. Mitra and J. B. Stark, Nature \textbf{411}, 1027 (2001).

%==================
\bibitem{Desurvire:2006}
E. Desurvire, J. Ligthwave Technol. \textbf{24}, 4697 (2006).

%==================
\bibitem{Essiambre:2008}
R.-J.~Essiambre, G.~J. Foschini, G. Kramer, and P.~J. Winzer, Phys. Rev.  Lett.
\textbf{101}, 163901 (2008).

%==================
\bibitem{Essiambre:2010}
R.-J.~Essiambre, G. Kramer, P.~J. Winzer, G.~J. Foschini, and B. Goebel,   J. of
Lightwave Technol.  \textbf{28}, 0733, (2010).

%==================
\bibitem{Ellis:2010}
A. D. Ellis, J. Zhao, and D. Cotter,  J. of Lightwave Technol.  \textbf{28}, 423, (2010).

%==================
\bibitem{Richardson:2010}
D. J. Richardson,  Science  \textbf{330}, 327, (2010).

%==================
\bibitem{Killey:2011}
R. Killey and C. Behrens, J. Mod. Opt.  \textbf{58}, 1, (2011).

%==================
\bibitem{N4}  A.~D. Ellis and J. Zhao, { \it Impact of Nonlinearities on Fiber Optic Communications}, Springer, New York, (2011).

%==================
\bibitem{Mecozzi:2012}
A. Mecozzi, R.-J.~Essiambre, J. of Lightwave Technol.  \textbf{30}, 2011, (2012).

%==================
\bibitem{Agrell:2012} E. Agrell, arXiv: 1108.0391v3.

%==================
\bibitem{Agrell:2013}
E. Agrell, "Nonlinear Fiber Capacity," presented at the Eur. Conf. Opt. Commun. LondonU.K., paper We.4.D.3 (2013).

%==================
\bibitem{Agrell:2014}
E Agrell, A Alvarado, G Durisi, M Karlsson, arXiv:1403.3339

%==================
\bibitem{tdyt03} K.S. Turitsyn, S.A. Derevyanko, I.V. Yurkevich, and S.K.~Turitsyn, \prl \textbf{91}, 203901 (2003).
%==================
\bibitem{Turitsyn:2012}
K. S. Turitsyn, K. K. Turitsyn, Opt. Lett.  \textbf{37}, 0146, (2012).

%==================
\bibitem{Sorokina2014} M. A. Sorokina   and S. K. Turitsyn,  Nat. Comm. {\bf 5},  3861 (2014)

\bibitem{Feynman} R.~P. Feynman, A.~R. Hibbs, {\it Quantim
 mechanics and path integrals}, McGraw-Hill Book Company, New York, (1965).

%==================
\bibitem{Terekhov:2014}
I.~S.~ Terekhov, S.~S. Vergeles, and S.~K. Turitsyn, Phys. Rev. Lett. {\bf 113},
230602 (2014).

%==================
\bibitem{Iannoe:1998}
E.~Iannoe, F.~Matera, A.~Mecozzi, and M.~Settembre, \emph{Nonlinear  Optical
Communication Networks}, John Wiley \& Sons, New York, (1998).

%==================
\bibitem{Turitsyn:2000}
S.K. Turitsyn, S.B. Medvedev, M.P. Fedoruk, and E.G. Turitsyna, Phys.  Rev. E
\textbf{61}, 3127 (2000).

%==================
\bibitem{SupplMat}
Supplementary Materials.


%==================
\bibitem{Shannon:1949}
C.~E. Shannon, \emph{Communication in the presence of noise}, Proc. Institute of Radio Engineers, vol. \textbf{37}, 1, (1949). \\
V.~A.~Kotelnikov, \emph{On the capacity of the ether and wire in
Telecommunications}, Proceedings of the First All-Union Congress
on Technical Reconstruction of Communication. Union Energy Committee. (1933), In Russian.

%==================nondispersive
\bibitem{TTKhR:2015} I.~S.~ Terekhov, A.~V. Reznichenko,  Ya.~A.~Kharkov, and S.~K. Turitsyn, arXiv: 1508.05774.

%==================small gamma
\bibitem{TTKhR:2016} I.~S.~ Terekhov, A.~V. Reznichenko, and S.~K. Turitsyn, arXiv: 1607.01184.



\end{thebibliography}

\documentclass[aps,pra,twocolumn,showpacs,superscriptaddress]{revtex4}
\usepackage[T2A]{fontenc}
\usepackage[cp1251]{inputenc}
\usepackage[english]{babel}
\usepackage{graphicx}
\usepackage{color}
\usepackage{bm}
\usepackage{amssymb}
\usepackage[intlimits]{amsmath}
\usepackage{amsmath}

\newcommand{\be}{\begin{equation}}
\newcommand{\ee}{\end{equation}}
\newcommand{\ben}{\begin{equation*}}
\newcommand{\een}{\end{equation*}}

\begin{document}

%\begin{titlepage}
\hskip 2cm
\date{\today}
\title{Mutual information in nonlinear communication channel.  Preliminary analytical results in large SNR and small nonlinearity limit}

\author{I.~S.~Terekhov}
\email[E-mail: ]{I.S.Terekhov@gmail.com} \affiliation{Budker Institute of  Nuclear
Physics of Siberian Branch Russian  Academy of Sciences, Novosibirsk, 630090 Russia}
%\affiliation{School of Physics, University of New South Wales, Sydney 2052,
%Australia}
\affiliation{Novosibirsk State University, Novosibirsk, 630090 Russia}
\author{A.~V.~Reznichenko}
\email[E-mail: ]{A.V.Reznichenko@inp.nsk.su} \affiliation{Budker Institute of
Nuclear Physics of Siberian Branch Russian Academy of Sciences, Novosibirsk, 630090
Russia} \affiliation{Novosibirsk State University, Novosibirsk, 630090 Russia}
\author{S.~K.~Turitsyn}
\email[E-mail: ]{s.k.turitsyn@aston.ac.uk} \affiliation{Novosibirsk State
University, Novosibirsk, 630090 Russia}
\affiliation{Aston Institute of Photonics
Technologies, Aston University, Aston Triangle, Birmingham, B4 7ET, UK}

\begin{abstract}

Applying perturbation theory to the path-integral representation for the mutual
information of the nonlinear communication channel described by the nonlinear
Shr\"{o}dinger equation (NLSE) with the additive Gaussian noise we analyze the
analytical expression for the mutual information at large signal-to-noise ratio
($\mathrm{SNR}$) and small nonlinearity. We classify all possible corrections to the
mutual information in nonlinearity parameter and demonstrate that all singular in
$\mathrm{SNR}$ terms vanish in the final result. Furthermore our analytical result
demonstrates that the corrections to Shannon's contribution to the mutual
information in the leading order in $\mathrm{SNR}$ are of order of squared
nonlinearity parameter. We outline the way for the calculation of these corrections
in the further investigations.

\end{abstract}

\keywords{information entropy, channel capacity, mutual information, nonlinear Shr\"{o}dinger equation.}
\pacs{05.10.Gg, 89.70.-a,  02.70.-c,02.70.Rr,05.90.+m}
\maketitle
%==================Introduction

\section{Introduction}

The channel capacity is one of the central concepts of information theory that has
its roots in statistical physics \cite{Shannon:1948}. The capacity  introduced by
Shannon in his seminal work \cite{Shannon:1948}  gives the maximum rate at which
information can be reliably transmitted through a noisy communication channel. The
channel capacity (in bits per symbol) is formally defined as a maximum of the mutual
information $I_{P[X]}$ over the input signal probability distribution functional
(PDF) $P[X]$:
\begin{eqnarray}
C=\max_{P[X]}  I_{P[X]}, \label{capacity1}
\end{eqnarray} under the condition of a fixed average power. The mutual information
$I_{P[X]}$  (in continuous-input, and continuous-output channel) is expressed
through the path-integral over input $X$ and output $Y$ signals:
\begin{eqnarray}
\begin{split} % I deleted \lim_{T\to\infty}\frac{1}{T}
&I_{P[X]}= \int {\cal D}X {\cal D}Y   P[X] P[Y| X]  \log\Big[ \frac{P[Y|
X]}{P_{out}[Y]}\Big], \label{capacity2}
\end{split}
\end{eqnarray}
where the output signal PDF $P_{out}[Y]$ reads
\begin{align}
& P_{out}[Y]=\int {\cal D}X P[X] P[Y|X],
\label{Pout}
\end{align}
with  $P[Y| X]$ being  the conditional probability density,   that is, the probability
of receiving  output signal $Y$ when the input signal is $X$. Both $X$ and $Y$ may be discrete or continuous. When $X$ is discrete,
notation integral over $X$ stands for the summation of an under integral function over its discrete support. Capacity in bits per symbol multiplied by the rate at which symbols are transmitted (in symbols per second) gives the error-free information transmission rate in bits per second.

The above definition (\ref{capacity2}) exemplifies that  channel capacity has a
close link to information entropy \cite{Shannon:1948}. Mutual information is a
difference between the entropy of the output signal $H[Y]=-\int{\cal D}Y P_{out}[Y]
\log\Big[P_{out}[Y]\Big] $ and conditional entropy $H[Y|X]=-\int {\cal D}X {\cal D}Y
P[X]P[Y| X] \log\Big[P[Y| X]\Big]$ (having a meaning of a measure of the uncertainty
about the output field $Y$ if the input field $X$ is known). When the signal and the
noise are independent variables and the received signal $Y$ is the sum of the
transmitted signal $X$ and the noise, then it can be shown explicitly that the
entropy is generated during transmission in noisy channel: $H[Y] \geq H[X]$. In this
case, the transmission rate is the entropy of the received signal less the entropy
of the noise. The maximum  of the functional (\ref{capacity1}), i.e. the channel
capacity, can be calculated for such linear channels with an additive white Gaussian
noise (AWGN):
\begin{eqnarray}\label{CapacityShannon}
C\propto\log\left(1+\mathrm{SNR}\right)\,,
\end{eqnarray}
where $\mathrm{SNR}$ is a signal-to-noise power ratio \cite{Shannon:1948}.

This seminal theoretical result is the foundation  of the communication theory and
it has proven its importance in a number of practical applications. To some extent,
the Eq. (\ref{CapacityShannon}) worked so well in so many situations that some
engineers cease to distinguish the general Shannon expression for capacity
(\ref{capacity1}) and particular result for the specific linear additive white
Gaussian noise  channel (\ref{CapacityShannon}). However, recent advances in
fibre-optic communication where the channel is nonlinear, as opposed to the linear
AWGN, changed the situation. To increase the channel capacity over a certain
bandwidth with a  given accumulated noise of optical amplifiers, one has to increase
the signal power, see (\ref{CapacityShannon}). This works in the low SNR limit but
the effect the refraction index's dependence on light intensity dramatically changes
the propagation properties of the channel at higher optical signal power. In other
words, the fibre-optic channel is nonlinear. Recent studies have shown that the
spectral efficiency (that is, the number of bits transmitted per second per Hertz
--- practical characteristics having the same dimension as channel capacity) of a
fibre-optic channel is limited by the Kerr nonlinearity. These studies indicated
that observable spectral efficiency always turns out to be less  than the Shannon
limit of the corresponding linear AWGN channel (\ref{CapacityShannon})
\cite{Mitra:2001,Desurvire:2006,Essiambre:2008,Essiambre:2010,Ellis:2010,Richardson:2010,Killey:2011,N4,Mecozzi:2012}.
It has been observed that  the spectral efficiency of the nonlinear channel
decreases with increasing $\mathrm{SNR}$ at high enough values of $\mathrm{SNR}$
\cite{Mitra:2001,Desurvire:2006,Essiambre:2008,Essiambre:2010,Ellis:2010,N4,Mecozzi:2012}.
This analysis certainly provides only a lower bound on channel capacity and  does
not prove that the Shannon nonlinear fibre channel capacity is decreasing with
power; see, for example,  discussions in
\cite{Agrell:2012,Agrell:2013,Agrell:2014,tdyt03}.

In general, there is a widely spread opinion that nonlinear channel capacity is
always less than the capacity of the corresponding  linear AWGN channel for equal
$\mathrm{SNR}$. However, in Ref. \cite{Turitsyn:2012,Sorokina2014} the authors note
that  nonlinearity can be either destructive or constructive. Moreover, in Ref.
\cite{Agrell:2012} it was proved that the capacity of certain nonlinear channels
could not decrease with $\mathrm{SNR}$. The capacity of nonlinear fibre channels is
still an open problem of great practical and fundamental importance.

In this work, we estimate analytically the first nonzero correction to the mutual
information of the channel described by the NLSE with additive Gaussian noise. We
then calculate in the perturbation theory (at large $\mathrm{SNR}$ and small
nonlinearity) the conditional probability density for the NLSE channel by the method
developed recently in \cite{Terekhov:2014}. Finally, we demonstrate that all
singular in $\mathrm{SNR}$ corrections  to the mutual information are cancelled and
the resulting correction is of order of squared nonlinearity parameter.

The article is organized as follows. In the next section we present the expression
for the conditional probability density. Then in the presented channel model we
calculate the mutual information and classify all possible correction to it at large
$\mathrm{SNR}$ and small nonlinearity parameter. Finally, we present the main
conclusions of the paper. Details of our calculations are placed in the
Supplementary Materials \cite{SupplMat}.

%==================The model of nonlinear channel and conditional probability
\section{The nonlinear channel model and conditional probability}

Let us consider the propagation of the signal  $\psi_\omega(z)$ in  the channel
modelled by the NLSE with AWGN, that we rewrite, for convenience, in the frequency
domain, see \cite{Iannoe:1998,Turitsyn:2000}:
\begin{eqnarray}\label{startingCannelEq}
&&\!\!\!\!\!\!\!\!\!\partial_z \psi_\omega (z)-i\beta\omega^2\psi_\omega (z) -
i\gamma \int \frac{d\omega_1 d\omega_2 d\omega_3}{(2\pi)^3} \times \nonumber\\
&& \!\!\!\!\!\!\!\!\! \times\delta(\omega+\omega_3-\omega_1-\omega_2)
\psi_{\omega_1} (z) \psi_{\omega_2} (z) \bar{\psi}_{\omega_3} (z)= \eta_\omega(z)\,,
\end{eqnarray}
where $\beta$ is the dispersion coefficient, $\gamma$ is the Kerr nonlinearity
coefficient, bar means complex conjugation, $\eta_\omega (z)$ is an additive complex
white noise with zero mean and the correlator $\langle \eta_\omega
(z)\bar{\eta}_{\omega^\prime} (z^\prime)\rangle_{\eta} = 2\pi Q
\delta(z-z^\prime)\delta(\omega-\omega^\prime)\, $, where $Q$ is  the noise power
per unit frequency per unit length \cite{Iannoe:1998,Essiambre:2010}:
$P_{noise}=QLW/2\pi$, here $W/(2 \pi)$ is a frequency bandwidth, and $L$ is the
channel length.

It is worth emphasizing that in a nonlinear channel, transmitted and received signal
bandwidths can be different. Therefore,  we assume here that in general, the input
$X(\omega)$  and output $Y(\omega)$ signals  may have different channel bandwidths
$W$ and $W'$, with $ W' \supset W$.

We introduce the dimensionless parameter $\tilde{\gamma}= P_{ave} \gamma L $ which
describes the impact of nonlinearity. Here the average power of the signal $X$ reads
\begin{align}\label{Pave}
P_{ave}=\lim_{T\to\infty}\int {\cal D}X P[X] \frac{1}{T}\int \frac{d\omega}{2\pi} |X({\omega})|^2\,,
\end{align}
where $T$ is the large time interval containing the whole input signal. It has been
shown in Ref. \cite{Terekhov:2014} that for NLSE channel governed by Eq.
(\ref{startingCannelEq}) in the large SNR limit,
\begin{eqnarray}
\epsilon =1/\mathrm{SNR}=QLW/(2\pi P_{ave})\ll 1,
\end{eqnarray}
the conditional  probability density $P[Y(\omega)| X(\omega)]$ to receive
$\psi_\omega(L)=Y(\omega)$  given the input signal $\psi_\omega(0)=X(\omega)$  can
be written as
\begin{widetext}
\begin{eqnarray}
P[Y(\omega)|X(\omega)]=\exp[-\frac{S[\Psi_{\omega}(z)]}{Q}\Big]
\int^{\tilde{\psi}_{\omega}(L)=0}_{\tilde{\psi}_{\omega}(0)=0} {\cal D} \tilde{\psi}
\exp\Big[
-\frac{1}{Q}\Big\{S[\Psi_{\omega}(z)+\tilde{\psi}_{\omega}(z)]-S[\Psi_{\omega}(z)]\Big\}\Big]\,,\label{PYXquasiclassic}
\\
S[\psi]=\int^L_0 dz \int \frac{d\omega}{2\pi} \Big|\partial_z \psi_{\omega}(z)-i
\beta \omega^2 \psi_{\omega}(z) - i\gamma \int \frac{d\omega_1 d\omega_2
d\omega_3}{(2\pi)^3} \delta(\omega+\omega_3-\omega_1-\omega_2) \psi_{\omega_1}(z)
\psi_{\omega_2}(z) \overline{\psi}_{\omega_3}(z) \Big|^2\,.\label{Action}
\end{eqnarray}
\end{widetext}
Here $\Psi_{\omega}(z)$ is  the so-called ``classical trajectory'' \cite{Feynman} of
the path-integral (\ref{PYXquasiclassic}), that is, the extremum function of the
action (\ref{Action}), i.e. $\delta S[\Psi]=0$~--- see Eq.~(6) in \cite{SupplMat},
with the boundary conditions $\Psi_{\omega}(0)=X(\omega)$,
$\Psi_{\omega}(L)=Y(\omega)$. At small $\tilde{\gamma}$ we can calculate
$P[Y(\omega)|X(\omega)]$, Eq. (\ref{PYXquasiclassic}), analytically using the
perturbation theory developed in Ref. \cite{Terekhov:2014}. There are two types of
terms in the expansion of Eq.~(\ref{PYXquasiclassic}) in $\tilde{\gamma}$. The first
type of perturbative corrections comes from the expansion of exponent
$\exp\left[-{S[\Psi_{\omega}(z)]}/{Q}\right]$ and has the structure
\begin{eqnarray}\label{ExponentExpansion}
&&\exp\left[-\frac{S[\Psi_{\omega}(z)]}{Q}\right]\approx\exp\left[-\frac{S[\Psi^{(0)}_{\omega}(z)]}{Q}\right]_{\gamma=0}\times\nonumber\\
&&\left(1+\sum_{p=0}^\infty\sum_{k=1}^\infty \alpha_{p,k}[\Psi^{(0)}_{\omega}(z)]\tilde{\gamma}^p\left(\frac{\tilde{\gamma}}{\epsilon}\right)^k\right)\,.
\end{eqnarray}
Here $\Psi^{(0)}_{\omega}(z)=e^{i\beta \omega^2 z}\Big[z B(\omega)/L+X(\omega)\Big]$
is the solution at $\gamma=0$ of the equation $\delta S[\Psi]=0$, see  Eqs.~(6) and
(8) in \cite{SupplMat}, with the boundary conditions $\Psi_{\omega}(0)=X(\omega)$,
$\Psi_{\omega}(L)=Y(\omega)$. Here $B(\omega) =e^{-i\beta \omega^2 L}
Y(\omega)-X(\omega) $, and $\alpha_{p,k}[\Psi^{(0)}_{\omega}(z)]$ are some
constructively defined functionals, see explicit expressions in \cite{SupplMat}.
The second type of corrections originates from the expansion of the path-integral in
Eq.~(\ref{PYXquasiclassic}) and has the form
\begin{eqnarray}\label{PathIntegralExpansion}
\Lambda^{(M^\prime)}_{QL}\left(1+\sum_{p=1}^\infty\sum_{k=0}^\infty\gamma_{p,k}
[\Psi^{(0)}_{\omega}(z)]\tilde{\gamma}^p\left(\tilde{\gamma}\epsilon\right)^k\right)\,,
\end{eqnarray}
where $\Lambda^{(M^\prime)}_{QL} = \Big( \frac{\delta}{\pi QL}\Big)^{M^\prime} $ is
the normalization factor and $\gamma_{p,k}[\Psi^{(0)}_{\omega}(z)]$ are some defined
functionals, see \cite{SupplMat}. Since the parameter $\epsilon$ is assumed to be
small, we can use the quasi-classical approach, and the main contribution to
$P[Y|X]$ comes from the expansion of exponent Eq. (\ref{ExponentExpansion}). We
substitute the function $\Psi^{(0)}$ to the right-hand side of Eq.
(\ref{ExponentExpansion}), then the obtained result is multiplied by
(\ref{PathIntegralExpansion}) leading to:
\begin{eqnarray}\label{fullExpansion}
&& P[Y(\omega)|X(\omega)]\approx P^{(0)}[Y(\omega)|X(\omega)]\Biggl(1+\gamma_{1,0}\tilde{\gamma}+\nonumber\\
&&\!\!\!\sum_{k=1}^\infty \alpha_{0,k}\left(\frac{\tilde{\gamma}}{\epsilon}\right)^k
+ \tilde{\gamma}\sum_{k=1}^\infty \left[\alpha_{1,k}+\alpha_{0,k}\gamma_{1,0}\right]
\left(\frac{\tilde{\gamma}}{\epsilon}\right)^k \Biggr)+\nonumber\\
&&{\cal O }\left( \tilde{\gamma}^2\right),
\end{eqnarray}
where
\begin{eqnarray}
 \!\! P^{(0)}[Y(\omega)|X(\omega)]=\Lambda^{(M^\prime)}_{QL}\exp\Big[- \!\frac{1}{QL}\! \!\int_{W'}\!
 \frac{d\omega}{2\pi} |B(\omega)|^2\Big].
\label{P0YX}
\end{eqnarray}
In Eq. (\ref{fullExpansion}) we keep only leading and next-to-leading  order  in
$\tilde{\gamma}$ terms for every order in  $\tilde{\gamma}/\epsilon$. This means
that the parameter $\tilde{\gamma}/\epsilon$ can be of  order of unity. We omit the
dependence of functionals  $\alpha_{p,k}$ and $\gamma_{p,k}$ on
$\Psi^{(0)}_{\omega}(z)$, but have it in mind. Now nonlinear corrections to the
mutual information can be calculated.

%=============Calculation of the nonlinear corrections to the mutual information
\section{ Calculation of the nonlinear corrections to the mutual information}

To calculate the mutual information, we have to calculate the path-integral
(\ref{capacity2}). In the previous section, we derived the expression
(\ref{fullExpansion}) for the conditional probability function.  Let us introduce a
probability function of input signal $P[X(\omega)]$.  In our consideration, the PDF
$P[X(\omega)]$  is chosen to be Gaussian in the spectral domain W and zero in $W'
\setminus W$. In discrete form  $P[X(\omega)]$ reads
\begin{eqnarray}
&\mspace{-20mu}P[X(\omega)]=  \Lambda^{(M)}_{P}\!\! \left(\prod^{M}_{i \in W}
e^{-\frac{\delta}{P} |X_i|^2}\right) \prod^{M'-M}_{j \in W' \setminus W}
\mspace{-2mu} \delta( X_{j}).\,\label{PX}
\end{eqnarray}
where $\delta(X_{j})=\delta( Re\, X_{j}) \delta(Im\, X_{j})$ is  the
$\delta$-function, frequency domain $W$ ($W'$) and is divided by $M$ ($M'$) grid
spacing $\delta = \frac{W}{2\pi M} = \frac{W'}{2\pi M'}$; $X_{j}=X(\omega_j)$. The
coefficient $\Lambda^{(M)}_{P}= \Big( \frac{\delta}{\pi P}\Big)^M $ follows from the
normalization condition $\int {\cal D}X P[X(\omega)] =1$. The measure ${\cal D}X$
 in (\ref{capacity2}), and in (\ref{Pout}), is understood as ${\cal D}X
=\prod_{j=1}^{M'} d Re\, X_j \, d Im \, X_j$.  As follows from the
Nyquist-Shannon-Kotelnikov theorem \cite{Shannon:1949} $M$ should be chosen greater
than $T {W}/{2\pi}$: the limit ${T\to\infty}$ in Eq.~(\ref{Pave}) is equivalent to
$M=T W/2\pi \to \infty$ in the measure. Parameter $P$ in Eq.~(\ref{PX}) describes
the signal power per unit of frequency (power spectral density), so the average
signal power (\ref{Pave}) is $P_{ave}=P W/2\pi \gg P_{noise}$ and the nonlinearity
parameter is $\tilde{\gamma}={\gamma} P L W/(2\pi)$.

First of all, corrections to the mutual information proportional to $\tilde{\gamma}$
vanish. These corrections to the mutual information come only from the term
$\gamma_{1,0}\tilde{\gamma}$ in (\ref{fullExpansion}). From the explicit expression
for $\gamma_{1,0}$  one can see that after integration over fields $X(\omega)$ and
$Y(\omega)$ these contributions vanish as the imaginary part of the real number, see
\cite{SupplMat}. Further, all corrections of order of $(\tilde{\gamma}/\epsilon)^k$
to the mutual information are equal to zero for all $k>0$ as well. We explain this
cancellation in \cite{SupplMat} by the counting of field $B(\omega)=e^{-i\beta
\omega^2 L} Y(\omega)-X(\omega)$ after the change of integration variable from
$Y(\omega)$ to $B(\omega)$ in the path-integral (\ref{capacity2}). Owing to
Eq.~(\ref{P0YX}) the variation scale of $B(\omega)$ in the path-integral is of order
of $\sqrt{QL}$. And every field $B(\omega)$ before the exponent yields the
suppression factor $\sqrt{QL}$ after the integration over $B(\omega)$. The first
next-to-leading in $\tilde{\gamma}$ singular correction  of order of
$\tilde{\gamma}^2/ \epsilon$ is equal to zero after the intricate cancellation of
different terms in the mutual information expression. We present the  details of
this cancellation  in \cite{SupplMat}. Basing on another approach to the mutual
information calculation \cite{TTKhR:2016} we can make a guess that all singular in
$1/\epsilon$ corrections (i.e. of order of
$\tilde{\gamma}^p(\tilde{\gamma}/\epsilon)^k$ for $k>0$ and $p\geq 0$) to the mutual
information should be equal to zero as well. Roughly speaking here we can use the
same arguments  as in the case of the leading singular corrections
$(\tilde{\gamma}/\epsilon)^k$. Finally, after all cancellations we obtain the
following expression for the mutual information:
\begin{eqnarray}
I_{P[X]}=M \log\Big[1+\frac{1}{\epsilon}\Big]+ {\cal O }\left(
\tilde{\gamma}^2\right)+{\cal O }\left( \epsilon \right)\,, \label{MutInfRes}
\end{eqnarray}
where the first term in the right-hand side is the well-known Shannon's  result
(\ref{CapacityShannon}) for the  linear channel \cite{Shannon:1948}, and the second
term is the nonlinear correction to the mutual information. As a coefficient of
${\cal O }\left( \tilde{\gamma}^2\right)$ one has the  function of dimensionless
dispersion parameter $\tilde{\beta}=\beta L W^2$. This function of $\tilde{\beta}$
should be negative at least for small dispersion parameter. Indeed for a
nondispersive nonlinear optical fibre channel one has the exact in nonlinearity
result \cite{TTKhR:2015} with the negative correction:
\begin{eqnarray}
&& \!\!\!\!I^{(\beta=0)}_{P[X]}=M \log\Big[1+\frac{1}{\epsilon}\Big]- \frac{M}{2} \!
\int^{\infty}_{0} \! \!d \tau e^{-\tau} \log\left(1+\frac{\tau^2
\tilde{\gamma}^2}{3} \right)= \nonumber
\\&& \!\!\!\! M \log\Big[1+\frac{1}{\epsilon}\Big]- M\frac{\widetilde{\gamma}^2}{3}+{\cal
O}(\widetilde{\gamma}^{\,4}). \label{MutInfzero}
\end{eqnarray}
Moreover at  large dispersion parameter $\tilde{\beta}$ the function in question
should be decreasing function approaching to the Shannon limit: for large dispersion
parameter one can neglect the nonlinear term in NLSE (\ref{startingCannelEq})
resulting in the linear channel, see Eq.~(\ref{CapacityShannon}). The exact
calculation of the $\tilde{\beta}$-dependence of the coefficient before ${\cal O
}\left( \tilde{\gamma}^2\right)$ in Eq.~(\ref{MutInfRes}) is the matter of our
further considerations \cite{TTKhR:2016}.

%==========================================Conclusion
\section{Conclusion}

We have described the perturbative method to obtain the analytical expression for
the mutual information $I_{P[X]}$ (\ref{capacity2}) of the NLSE channel at large
$\mathrm{SNR}={1}/{\epsilon}$ and small nonlinearity $\tilde{\gamma}\ll 1$. We have
demonstrated that all singular in $\epsilon$ terms in the leading orders
$(\tilde{\gamma}/\epsilon)^k$ and in the first next-to-leading order
$\tilde{\gamma}^2/ \epsilon$ vanish in the expression for the mutual information
$I_{P[X]}$ calculated for the Gaussian PDF $P[X]$. At small nonlinearity
$\tilde{\gamma}$ the first nonlinear correction to the mutual information is of
order of $\tilde{\gamma}^2$, and it is negative at least for small dispersion
parameter $\tilde{\beta}$ and vanishing for large $\tilde{\beta}$. Let us stress
once again that we have considered only the mutual information $I_{P[X]}$ in the
case of a Gaussian input signal PDF $P[X]$ rather than the channel capacity
(\ref{capacity1}). However, the quantity $I_{P[X]}$ is a natural estimate of a low
bound on the channel capacity since it reproduces the Shannon capacity of the linear
AWGN channel (\ref{CapacityShannon}) in the leading order in $\tilde{\gamma}$. The
ultimate calculation of the corrections of order of $\tilde{\gamma}^2$ to the mutual
information will be the matter of our future considerations, see \cite{TTKhR:2016}.

%==========================================Acknowledgments
\emph{\it Acknowledgments.}

The work of I.~S.~Terekhov was supported by the Russian Science  Foundation (RSF)
(grant No. 16-11-10133). Part of the work was supported by the Russian Foundation
for Basic Research (RFBR), Grant No. 16-31-60031/15. A.~V.~Reznichenko thanks the
President program (СП-2415.2015.2) for support. The work of S.~K.~Turitsyn was
supported by the grant of the Ministry  of Education and Science of the Russian
Federation (agreement No. 14.B25.31.0003) and the EPSRC project UNLOC.

%===========================================Bibliography

\end{document}